OpenStudy (anonymous):

@MJgirL1958 Last month, Gavin had $5460.40 that he divided into two bank accounts. He deposited 3/8 into a savings account. He deposited the rest into his checking account. How much money did Gavin deposit into his checking account? I also fan and medal

OpenStudy (anonymous):

One way to do it is to divide $5460.40 by 8 so we will find out how much is 1/8 of that sum: 5460.40 / 8 = 682,55 Now 3/8 of the sum is 682,55 * 3 = 2047,65. That's how much Gavin deposited into a savings account. 5460,40 - 2047,65 = 3412,75. That's 5/8 of his money which went to his checking account.

OpenStudy (anonymous):

so is 3412,75

OpenStudy (texaschic101):

3/8 of 5460.40 = 3/8 * 5460.40 = 16381.20/8 = 2047.65 that is what was deposited in savings account. 5460.40 - 2047.65 = 3412.75 yes...that is the answer

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

no problem :)